OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of the fuction y=-x^2+6x-11? (3,-2) (3,16) (-3,-29) (-3,-20)

OpenStudy (anonymous):

first coordinate is always \(-\frac{b}{2a}\) which in your case is \[-\frac{6}{2\times (-1)}=3\]

OpenStudy (anonymous):

second coordinate is what you get when you replace \(x\) by the first coordinate